Is the line through (3,-5) and (-1,7) and the line through (6,-13) and (-2,11) parallel, pependicular, or neither?
Answer by unlockmath(1688)

Hello,
Use the formula to find the slope of a line (y-y/x-x)
(-5-7/3--3)=-12/6=-2
(-13-11/6--2)=-24/8=-3
Since they not the same slope it's NOT parallel and they are not the negative inverse so it is NOT perpendicular so the answer is neither.
Make sense?
RJ
www.math-unlock.com